OpenStudy (anonymous):

A system of equations is shown below: 2x = −y + 6 −4x + 3y = 8 What is the solution to this system of equations? (−1, −4) (1, 4) (4, 1) (−4, −1)

pooja195 (pooja195):

2x = −y + 6 <---equation 1 −4x + 3y = 8 <---equation 2 Since @loverlarap mentioned substitution we'll use it :) Start by isolating a vairble like this: Solve 2x=−y for y: 2x+y=−y+6+y(Add y to both sides) 2x+y=6 2x+y+−2x=6+−2x(Add -2x to both sides) y=−2x+6 Now you have one new equation +the old one y=-2x+6 <---equation 1 −4x + 3y = 8 <----equation 2 Take the y equals part and plug it into equation 2 as we are substituting -4x+3(-2x+6)=8 solve for x

OpenStudy (anonymous):

Try substituting your answers in to the equations and see if it is equal.

OpenStudy (anonymous):

sorry for the late reply @pooja195 @Lovelarap -4x+3(-2x+6)=8 ..... x = 1

OpenStudy (anonymous):

yes,continue.

OpenStudy (anonymous):

what was the other equation? i got a little confused with all the words lol

OpenStudy (anonymous):

substitute x=1 in to one of the equations and you should get y.